1/1

Mu Du,  Mu Du,
Jiāngnán Chūn  Jiangnan Spring

translated by N. N. 72

Qiān lǐ yīng tí lù yìng hóng Shuǐ cūn shān guō jiǔ qí fēng. Nán cháo sì bǎi bā shí sì Duōshao lóu tái yān yǔ zhōng.

A thousand miles of oriole songs and red among the green Of wine flags flapping along the shore and in the hills Four hundred and eighty temples built by the Southern Court And how many pagodas in the land of mist and rain.

Author of this ĉina poem is Mu Du (Du Mu, ∗803 – †852).

Arg-938-1880 (2013-11-22 19:14:52)

Vidu ankaŭ: http: // www. mountainsongs. net/ poem_ .php? id= 839 .

Translation of the ĉina poem “Jiāngnán Chūn” by Mu Du (Du Mu, ∗803 – †852) into English by N. N. 72.

Arg-938-1885 (2013-11-22 19:44:06)

Tiu ĉi laŭvorta Angligo de la poemo troviĝas en http: // www. mountainsongs. net/ poem_.php? id= 839 .

http://www.poezio.net/version?poem-id=938&version-id=1880,1885